A 325 kg uniform square sign, 2/00 m on a side, is
Answer #1

length of the cable = r= sqrt(y^2+x^2)= sqrt(4^2+3^2) = 5 m

\let theta be the angle made with vertical

tantheta = x/y


theta = 37

In equilibrium net torque = 0

torque due to cable = T*cos37*x


torque due to sign board = -325*2


net torque = 0

T*cos37*x - 325*2 = 0

T*cos37*3 - 325*2 = 0


T = 271.3 N

+++++++++++++++++++++++++


along horizantal Fnet = 0

T*sin37 = Fh


Fh = 163.3 N


______________


along vertical


Fnet = 0

T*cos37 = mg

271.3*cos37 + Fv = 325*9.8


Fv = 2968.3 N